OpenStudy (anonymous):

20% of what number is 47

OpenStudy (anonymous):

47*5=235

OpenStudy (anonymous):

20x/100=47 You just asked a similar question. OpenStudy does not support those who are only looking for answers and not doing their own homework.

OpenStudy (anonymous):

Because 20% = 1/5 find the reciprocal

Parth (parthkohli):

Let's say that it is 20% of x. Equate. \(\Large \color{Black}{\Rightarrow {20 \over 100}x = 47}\) \(\Large \color{Black}{\Rightarrow {x \over 5} = 47 }\) \(\Large \color{Black}{\Rightarrow x = 47 \times 5 = ? }\)

OpenStudy (anonymous):

i am doing my homework im just still little confused. gosh

Parth (parthkohli):

What is 47 times 5?

OpenStudy (anonymous):

235

Parth (parthkohli):

That's it :D